A driver accelerates a 243-kg snowmobile, which results in a force being exerted that speeds the

snowmobile up from 7 m/s to 34 m/s over a time interval of 60 s. What is the magnitude of the average force that is exerted on the snowmobile?


1
Expert's answer
2021-03-29T09:58:40-0400

Let's first find the acceleration of the snowmobile:


a=vv0t=34 ms7 ms60 s=0.45 ms2.a=\dfrac{v-v_0}{t}=\dfrac{34\ \dfrac{m}{s}-7\ \dfrac{m}{s}}{60\ s}=0.45\ \dfrac{m}{s^2}.

Finally, we can find the magnitude of the average force that is exerted on the snowmobile:


F=ma=243 kg0.45 ms2=109.35 N.F=ma=243\ kg\cdot0.45\ \dfrac{m}{s^2}=109.35\ N.